# Recursive construction and enumeration of self-orthogonal and self-dual codes over finite commutative chain rings of even characteristic - I

### v1

@@ 2 Some preliminaries
Moreover, the matrix $\mathtt{V}^{(\kappa-1)}]_{s-\kappa_{1}}$ satisfying (3.45)-(3.47) has precisely
$(2^{m})^{\sum\limits_{i=\kappa+1}^{s+\kappa_{1}+1}\lambda_{i}\Lambda_{i-\kappa}+\Lambda_{s-\kappa_{1}}(n-\Lambda_{s+\kappa_{1}+1})-2\Lambda_{s-\kappa+1}-\frac{\Lambda_{s-\kappa_{1}}(\Lambda_{s-\kappa_{1}}-1)}{2}+\omega_{\kappa+1}}$
distinct choices, where $\omega_{\kappa+1}=1$ if $\textbf{1}\in\mathcal{C}^{(s-\kappa+1)},$ while $\omega_{\kappa+1}=0$ otherwise. Further, for a given choice of the matrix $\mathtt{V}^{(\kappa-1)}]_{s-\kappa_{1}}$ satisfying (3.45)-(3.47), we can write
@@
$2^{\epsilon}(2^{m})^{\sum\limits_{j=\kappa+2}^{s+\kappa_{1}+1}\lambda_{j}\Lambda_{j-\kappa-1}+\Lambda_{s-\kappa_{1}}(n-\Lambda_{s+\kappa_{1}+1})-\Lambda_{s-\kappa}-\frac{\Lambda_{s-\kappa_{1}}(\Lambda_{s-\kappa_{1}}-1)}{2}}$
@@
The following proposition lifts a self-orthogonal code of type $\{\Lambda_{s-\kappa_{1}+1},\lambda_{s-\kappa_{1}+2},\ldots,\lambda_{s+\kappa_{1}+1}\}$ and length $n$ over $\mathscr{R}_{\kappa,m}$ satisfying property $(\mathfrak{P})$ to a self-orthogonal code of type $\{\Lambda_{s-\kappa_{1}},\lambda_{s-\kappa_{1}+1},\ldots,\lambda_{s+\kappa_{1}+2}\}$ and the same length $n$ over $\mathscr{R}_{\kappa+2,m}$ satisfying property $(\mathfrak{P})$ when $e$ is odd and $2\kappa\leq e.$ It also enumerates the number of possible ways to perform this lifting.
@@
Let $e$ be odd, $2\kappa\leq e,$ and let $\ell=\kappa+2.$ Let $\mathscr{D}_{\kappa}$ be a self-orthogonal code of type $\{\Lambda_{s-\kappa_{1}+1},\lambda_{s-\kappa_{1}+2},\\ \ldots,\lambda_{s+\kappa_{1}+1}\}$ and length $n$ over $\mathscr{R}_{\kappa,m}$ satisfying property $(\mathfrak{P}).$ Let $\mathcal{C}^{(s-\kappa_{1})}$ and $\mathcal{C}^{(s-\kappa+1)}$ be linear subcodes of $Tor_{1}(\mathscr{D}_{\kappa}),$ such that $\mathcal{C}^{(s-\kappa+1)}\subseteq\mathcal{C}^{(s-\kappa_{1})},$ $\dim\mathcal{C}^{(s-\kappa+1)}=\Lambda_{s-\kappa+1},$ $\dim\mathcal{C}^{(s-\kappa_{1})}=\Lambda_{s-\kappa_{1}},$ and the code $\mathcal{C}^{(s-\kappa_{1})}$ is doubly even, (such a pair of codes $\mathcal{C}^{(s-\kappa_{1})}$ and $\mathcal{C}^{(s-\kappa+1)}$ exists for each choice of $\mathscr{D}_{\kappa},$ by Lemma 3.2). The following hold.
  1. (a)
    There exists a self-orthogonal code $\mathscr{D}_{\kappa+2}$ of type $\{\Lambda_{s-\kappa_{1}},\lambda_{s-\kappa_{1}+1},\ldots,\lambda_{s+\kappa_{1}+2}\}$ and length $n$ over $\mathscr{R}_{\kappa+2,m}$ satisfying property $(\mathfrak{P})$ with $Tor_{1}(\mathscr{D}_{\kappa+2})=\mathcal{C}^{(s-\kappa_{1})}$ and $Tor_{j+1}(\mathscr{D}_{\kappa+2})=Tor_{j}(\mathscr{D}_{\kappa})$ for $1\leq j\leq\kappa,$ where such a code exists:
    • unconditionally, if $\textbf{1}\notin\mathcal{C}^{(s-\kappa+1)};$
    • if and only if either $n\equiv 0\pmod{8}$ or $n\equiv 4\pmod{8}$ with $m$ even when $\textbf{1}\in\mathcal{C}^{(s-\kappa+1)}.$
  2. (b)
    Moreover, each such choice of the codes $\mathscr{D}_{\kappa},$ $\mathcal{C}^{(s-\kappa_{1})},$ and $\mathcal{C}^{(s-\kappa+1)}$ gives rise to precisely
    $2^{\epsilon}(2^{m})^{\sum\limits_{i=\kappa+2}^{s+\kappa_{1}+2}\lambda_{i}\Lambda_{i-\kappa-1}+\sum\limits_{j=\kappa+3}^{s+\kappa_{1}+2}\lambda_{j}\Lambda_{j-\kappa-2}+Y_{\kappa+2}}{\lambda_{s+\kappa_{1}+2}+n-\Lambda_{s+\kappa_{1}+2}-\Lambda_{s-\kappa_{1}}\brack\lambda_{s+\kappa_{1}+2}}_{2^{m}}$
    distinct self-orthogonal codes $\mathscr{D}_{\kappa+2}$ of type $\{\Lambda_{s-\kappa_{1}},\lambda_{s-\kappa_{1}+1},\ldots,\lambda_{s+\kappa_{1}+2}\}$ and length $n$ over $\mathscr{R}_{\kappa+2,m}$ satisfying property $(\mathfrak{P})$ with $Tor_{1}(\mathscr{D}_{\kappa+2})=\mathcal{C}^{(s-\kappa_{1})}$ and $Tor_{j+1}(\mathscr{D}_{\kappa+2})=Tor_{j}(\mathscr{D}_{\kappa})$ for $1\leq j\leq\kappa,$ where $Y_{\kappa+2}=(\Lambda_{s+\kappa_{1}+1}+\Lambda_{s-\kappa_{1}})(n-\Lambda_{s+\kappa_{1}+2}-\Lambda_{s-\kappa_{1}})+\Lambda_{s-\kappa_{1}}^{2}+\Lambda_{s-\kappa_{1}}-\Lambda_{s-\kappa+1}-\Lambda_{s-\kappa}$ and $\epsilon=1$ if $\textbf{1}\in\mathcal{C}^{(s-\kappa+1)},$ while $\epsilon=0$ otherwise.
Proof.
Working as in Proposition 3.4, the desired result follows.∎
The following proposition considers the case $2\kappa\leq e$ and provides a method for lifting a self-orthogonal code of type $\{\Lambda_{\gamma_{\ell}+2},\lambda_{\gamma_{\ell}+3},\ldots,\lambda_{\gamma_{\ell}+\ell-1}\}$ and length $n$ over $\mathscr{R}_{\ell-2,m}$ satisfying property $(\mathfrak{P})$ to a self-orthogonal code of type $\{\Lambda_{\gamma_{\ell}+1},\lambda_{\gamma_{\ell}+2},\ldots,\lambda_{\gamma_{\ell}+\ell}\}$ and the same length $n$ over $\mathscr{R}_{\ell,m}$ satisfying property $(\mathfrak{P}),$ where the integer $\ell$ satisfies $\kappa+3\leq\ell\leq e-\kappa+1$ and $\ell\equiv e\pmod{2}$. It also enumerates the number of possible ways to perform this lifting.
Proposition 3.6.
Let $2\kappa\leq e,$ and let $\ell$ be a fixed integer satisfying $\kappa+3\leq\ell\leq e-\kappa+1$ and $\ell\equiv e\pmod{2}$. Let us define $\gamma_{\ell}=s-\mathrm{f}_{\ell}.$ Let $\mathscr{D}_{\ell-2}$ be a self-orthogonal code of type $\{\Lambda_{\gamma_{\ell}+2},\lambda_{\gamma_{\ell}+3},\ldots,\lambda_{\gamma_{\ell}+\ell-1}\}$ and length $n$ over $\mathscr{R}_{\ell-2,m}$ satisfying property $(\mathfrak{P}).$ Let $\mathcal{C}^{(\gamma_{\ell}+1)}$ be a $\Lambda_{\gamma_{\ell}+1}$-dimensional linear subcode of the torsion code $Tor_{1}(\mathscr{D}_{\ell-2}).$ The following hold.
  1. (a)
  2. (b)
    Moreover, each such pair of codes $\mathscr{D}_{\ell-2}$ and $\mathcal{C}^{(\gamma_{\ell}+1)})$ gives rise to precisely
    $(2^{m})^{\sum\limits_{i=\ell}^{\gamma_{\ell}+\ell}\lambda_{i}\Lambda_{i-\ell+1}+\sum\limits_{j=\ell+1}^{\gamma_{\ell}+\ell}\lambda_{j}\Lambda_{j-\ell}+Y_{\ell}}{\lambda_{\gamma_{\ell}+\ell}+n-\Lambda_{\gamma_{\ell}+\ell}-\Lambda_{\gamma_{\ell}+1}\brack\lambda_{\gamma_{\ell}+\ell}}_{2^{m}}\vskip-5.69054pt$
    distinct self-orthogonal codes $\mathscr{D}_{\ell}$ of type $\{\Lambda_{\gamma_{\ell}+1},\lambda_{\gamma_{\ell}+2},\ldots,\lambda_{\gamma_{\ell}+\ell}\}$ and length $n$ over $\mathscr{R}_{\ell,m}$ satisfying property $(\mathfrak{P})$ with $Tor_{1}(\mathscr{D}_{\ell})=\mathcal{C}^{(\gamma_{\ell}+1)}$ and $Tor_{i+1}(\mathscr{D}_{\ell})=Tor_{i}(\mathscr{D}_{\ell-2})$ for $1\leq i\leq\ell-2,$ where $Y_{\ell}=(\Lambda_{\gamma_{\ell}+\ell-1}+\Lambda_{\gamma_{\ell}+1})(n-\Lambda_{\gamma_{\ell}+\ell}-\Lambda_{\gamma_{\ell}+1})+\Lambda_{\gamma_{\ell}+1}^{2}+\Lambda_{\gamma_{\ell}+1}-\Lambda_{\gamma_{\ell}+1-\kappa_{1}}-\Lambda_{\gamma_{\ell}-\kappa_{1}}.$
Proof.
Working as in Proposition 4.3 of Yadav and Sharma [31], the desired result follows immediately. ∎
The following proposition considers the case $2\kappa>e$ and provides a method for lifting a self-orthogonal code of type $\{\Lambda_{\gamma_{\ell}+2},\lambda_{\gamma_{\ell}+3},\ldots,\lambda_{\gamma_{\ell}+\ell-1}\}$ and length $n$ over $\mathscr{R}_{\ell-2,m}$ satisfying property $(\mathfrak{P})$ to a self-orthogonal code of type $\{\Lambda_{\gamma_{\ell}+1},\lambda_{\gamma_{\ell}+2},\ldots,\lambda_{\gamma_{\ell}+\ell}\}$ and length $n$ over $\mathscr{R}_{\ell,m}$ satisfying property $(\mathfrak{P}),$ where the integer $\ell$ satisfies $e-\kappa+2-2\theta_{e}\leq\ell\leq\kappa-\mathrm{f}_{2\kappa-e}+1$ and $\ell\equiv e\pmod{2}$. It also enumerates the number of possible ways to perform this lifting.
Proposition 3.7.
@@
Let $2\kappa>e,$ and let $\ell$ be a fixed integer satisfying $e-\kappa+2-2\theta_{e}\leq\ell\leq\kappa-\mathrm{f}_{2\kappa-e}+1$ and $\ell\equiv e\pmod{2}$. Let us define $\gamma_{\ell}=s-\mathrm{f}_{\ell}.$ Let $\mathscr{D}_{\ell-2}$ be a self-orthogonal code of type $\{\Lambda_{\gamma_{\ell}+2},\lambda_{\gamma_{\ell}+3},\ldots,\lambda_{\gamma_{\ell}+\ell-1}\}$ and length $n$ over $\mathscr{R}_{\ell-2,m}$ satisfying property $(\mathfrak{P}).$ Let $\mathcal{C}^{(\gamma_{\ell}+1)}$ be a $\Lambda_{\gamma_{\ell}+1}$-dimensional linear subcode of the torsion code $Tor_{1}(\mathscr{D}_{\ell-2}).$ The following hold.
@@
    $\displaystyle(2^{m})^{\sum\limits_{i=\ell}^{\gamma_{\ell}+\ell}\lambda_{i}\Lambda_{i-\ell+1}+\sum\limits_{j=\ell+1}^{\gamma_{\ell}+\ell}\lambda_{j}\Lambda_{j-\ell}+Y_{\ell}}{\lambda_{\gamma_{\ell}+\ell}+n-\Lambda_{\gamma_{\ell}+\ell}-\Lambda_{\gamma_{\ell}+1}\brack\lambda_{\gamma_{\ell}+\ell}}_{2^{m}}\vskip-5.69054pt$
    distinct self-orthogonal codes $\mathscr{D}_{\ell}$ of type $\{\Lambda_{\gamma_{\ell}+1},\lambda_{\gamma_{\ell}+2},\ldots,\lambda_{\gamma_{\ell}+\ell}\}$ and length $n$ over $\mathscr{R}_{\ell,m}$ satisfying property $(\mathfrak{P})$ with $Tor_{1}(\mathscr{D}_{\ell})=\mathcal{C}^{(\gamma_{\ell}+1)}$ and $Tor_{i+1}(\mathscr{D}_{\ell})=Tor_{i}(\mathscr{D}_{\ell-2})$ for $1\leq i\leq\ell-2,$ where $Y_{\ell}=(\Lambda_{\gamma_{\ell}+\ell-1}+\Lambda_{\gamma_{\ell}+1})(n-\Lambda_{\gamma_{\ell}+\ell}-\Lambda_{\gamma_{\ell}+1})+\Lambda_{\gamma_{\ell}+1}^{2}+\Lambda_{\gamma_{\ell}+1}-\Lambda_{s-2\mathrm{f}_{\ell}+2}-\Lambda_{s-2\mathrm{f}_{\ell}+1}.$
@@
By applying Lemma 4.1 of Yadav and Sharma [32] and working as in Proposition 3.3, the desired result follows immediately. ∎
The following proposition lifts a self-orthogonal code of type $\{\Lambda_{\gamma_{\ell}+2},\lambda_{\gamma_{\ell}+3},\ldots,\lambda_{\gamma_{\ell}+\ell-1}\}$ and length $n$ over $\mathscr{R}_{\ell-2,m}$ satisfying property $(\mathfrak{P})$ to a self-orthogonal code of type $\{\Lambda_{\gamma_{\ell}+1},\lambda_{\gamma_{\ell}+2},\ldots,\lambda_{\gamma_{\ell}+\ell}\}$ and the same length $n$ over $\mathscr{R}_{\ell,m}$ satisfying property $(\mathfrak{P}),$ where the integer $\ell$ satisfies $\ell\equiv e\pmod{2}$ and $e-\kappa+2\leq\ell\leq e$ if $2\kappa\leq e$, while $\kappa-\mathrm{f}_{2\kappa-e}+1<\ell\leq e$ if $2\kappa>e.$ It also enumerates the number of possible ways to perform this lifting.
Proposition 3.8.
Let $\ell$ be a fixed integer satisfying $\ell\equiv e\pmod{2}$ and $e-\kappa+2\leq\ell\leq e$ if $2\kappa\leq e$, while $\kappa-\mathrm{f}_{2\kappa-e}+1<\ell\leq e$ if $2\kappa>e.$ Let us define $\gamma_{\ell}=s-\mathrm{f}_{\ell}.$ Let $\mathscr{D}_{\ell-2}$ be a self-orthogonal code of type $\{\Lambda_{\gamma_{\ell}+2},\lambda_{\gamma_{\ell}+3},\ldots,\lambda_{\gamma_{\ell}+\ell-1}\}$ and length $n$ over $\mathscr{R}_{\ell-2,m}$ satisfying property $(\mathfrak{P}).$ Let $\mathcal{C}^{(\gamma_{\ell}+1)}$ be a $\Lambda_{\gamma_{\ell}+1}$-dimensional linear subcode of the torsion code $Tor_{1}(\mathscr{D}_{\ell-2}).$ The following hold.
  1. (a)
  2. (b)
    $\displaystyle(2^{m})^{\sum\limits_{i=\ell}^{\gamma_{\ell}+\ell}\lambda_{i}\Lambda_{i-\ell+1}+\sum\limits_{j=\ell+1}^{\gamma_{\ell}+\ell}\lambda_{j}\Lambda_{j-\ell}+Y_{\ell}}{\lambda_{\gamma_{\ell}+\ell}+n-\Lambda_{\gamma_{\ell}+\ell}-\Lambda_{\gamma_{\ell}+1}\brack\lambda_{\gamma_{\ell}+\ell}}_{2^{m}}\vskip-5.69054pt$
    distinct self-orthogonal codes $\mathscr{D}_{\ell}$ of type $\{\Lambda_{\gamma_{\ell}+1},\lambda_{\gamma_{\ell}+2},\ldots,\lambda_{\gamma_{\ell}+\ell}\}$ and length $n$ over $\mathscr{R}_{\ell,m}$ satisfying property $(\mathfrak{P})$ with $Tor_{1}(\mathscr{D}_{\ell})=\mathcal{C}^{(\gamma_{\ell}+1)}$ and $Tor_{i+1}(\mathscr{D}_{\ell})=Tor_{i}(\mathscr{D}_{\ell-2})$ for $1\leq i\leq\ell-2,$ where $Y_{\ell}=(\Lambda_{\gamma_{\ell}+\ell-1}+\Lambda_{\gamma_{\ell}+1})(n-\Lambda_{\gamma_{\ell}+\ell}-\Lambda_{\gamma_{\ell}+1})+\Lambda_{\gamma_{\ell}+1}^{2}+\Lambda_{\gamma_{\ell}+1}.$
Proof.
Working as in Proposition 4.3 of Yadav and Sharma [31], the desired result follows immediately. ∎
@@
We next assert that $\textbf{1}\notin Tor_{s-\kappa+\theta_{e}}(\mathscr{D}_{e})$ when $2\kappa\leq e,$ $n\equiv 4\pmod{8}$ and $m$ is odd. To prove this assertion, let us suppose that the code $\mathscr{D}_{e}$ has a generator matrix $\mathcal{G}_{e}$ as defined in (3.1). Further, corresponding to the code $\mathscr{D}_{e},$ let us consider a linear code $\mathscr{D}_{\ell}$ of type $\{\Lambda_{\gamma_{\ell}+1},\lambda_{\gamma_{\ell}+2},\ldots,\lambda_{\gamma_{\ell}+\ell}\}$ and length $n$ over $\mathscr{R}_{\ell,m}$ with a generator matrix $\mathcal{G}_{\ell}$ (as defined in (3.2)) for each integer $\ell$ satisfying $2\leq\ell\leq e-2$ and $\ell\equiv e\pmod{2}$. By Lemma 3.1, we see that $\mathscr{D}_{\ell}$ is a self-orthogonal code over $\mathscr{R}_{\ell,m}$ satisfying property $(\mathfrak{P}).$ In particular, $\mathscr{D}_{\kappa-1+\theta_{e}}$ is a self-orthogonal code of type $\{\Lambda_{s-\kappa_{1}+1},\lambda_{s-\kappa_{1}+2},\ldots,\lambda_{s+\kappa_{1}+\theta_{e}}\}$ over $\mathscr{R}_{\kappa-1+\theta_{e},m}$ satisfying property $(\mathfrak{P})$ with $\mathscr{D}_{\kappa+1+\theta_{e}}$ is a self-orthogonal code of type $\{\Lambda_{s-\kappa_{1}},\lambda_{s-\kappa_{1}+1},\ldots,\lambda_{s+\kappa_{1}+1+\theta_{e}}\}$ over $\mathscr{R}_{\kappa+1+\theta_{e},m}$ satisfying property $(\mathfrak{P}).$ Now, working as in Propositions 3.4 and 3.5, it is easy to see that $\textbf{1}\notin Tor_{s-\kappa+\theta_{e}}(\mathscr{D}_{e})$ when $2\kappa\leq e,$ $n\equiv 4\pmod{8}$ and $m$ is odd. On taking $\mathcal{C}^{(i)}=Tor_{i}(\mathscr{D}_{e})$ for $1\leq i\leq s+\theta_{e},$ the desired result follows. ∎

### v2

@@ 2.6 Notations and preliminary setup
    Moreover, the matrix $\mathtt{V}^{(\kappa-1)}]_{s-\kappa_{1}}$ satisfying (3.45)–(3.47) has precisely
    $\vskip-2.84526pt\displaystyle(2^{m})^{\sum\limits_{i=\kappa+1}^{s+\kappa_{1}+1}\lambda_{i}\Lambda_{i-\kappa}+\Lambda_{s-\kappa_{1}}(n-\Lambda_{s+\kappa_{1}+1})-2\Lambda_{s-\kappa+1}-\frac{\Lambda_{s-\kappa_{1}}(\Lambda_{s-\kappa_{1}}-1)}{2}+\omega_{\kappa+1}}$
    distinct choices, where $\omega_{\kappa+1}=1$ if $\textbf{1}\in\mathcal{C}^{(s-\kappa+1)},$ while $\omega_{\kappa+1}=0$ otherwise. Further, for a given choice of the matrix $\mathtt{V}^{(\kappa-1)}]_{s-\kappa_{1}}$ satisfying (3.45)–(3.47), we can write
@@
    $\vskip-2.84526pt\displaystyle 2^{\epsilon}(2^{m})^{\sum\limits_{j=\kappa+2}^{s+\kappa_{1}+1}\lambda_{j}\Lambda_{j-\kappa-1}+\Lambda_{s-\kappa_{1}}(n-\Lambda_{s+\kappa_{1}+1})-\Lambda_{s-\kappa}-\frac{\Lambda_{s-\kappa_{1}}(\Lambda_{s-\kappa_{1}}-1)}{2}}$
@@
  • (II)
    When $e$ is odd, the desired result follows by working as in case (I).
The following proposition lifts a self-orthogonal code of type $\{\Lambda_{\gamma_{\ell}+2},\lambda_{\gamma_{\ell}+3},\ldots,\lambda_{\gamma_{\ell}+\ell-1}\}$ and length $n$ over $\mathscr{R}_{\ell-2,m}$ satisfying property $(\mathfrak{P})$ to a self-orthogonal code of type $\{\Lambda_{\gamma_{\ell}+1},\lambda_{\gamma_{\ell}+2},\ldots,\lambda_{\gamma_{\ell}+\ell}\}$ and the same length $n$ over $\mathscr{R}_{\ell,m}$ that also satisfies property $(\mathfrak{P}),$ where the integer $\ell$ satisfies $\ell\equiv e\pmod{2}$ and $\kappa+3\leq\ell\leq e$ when $2\kappa\leq e$, while it satisfies $\ell\equiv e\pmod{2}$ and $e-\kappa+2-2\theta_{e}\leq\ell\leq e$ when $2\kappa>e.$ Additionally, it determines the number of distinct ways in which this lifting can be performed.
@@
Let $\ell$ be a fixed integer satisfying $\ell\equiv e\pmod{2}$ and $\kappa+3\leq\ell\leq e$ if $2\kappa\leq e$, while the integer $\ell$ satisfies $\ell\equiv e\pmod{2}$ and $e-\kappa+2-2\theta_{e}\leq\ell\leq e$ if $2\kappa>e.$ Let us define $\gamma_{\ell}=s-\mathrm{f}_{\ell}.$ Let $\mathscr{D}_{\ell-2}$ be a self-orthogonal code of type $\{\Lambda_{\gamma_{\ell}+2},\lambda_{\gamma_{\ell}+3},\ldots,\lambda_{\gamma_{\ell}+\ell-1}\}$ and length $n$ over $\mathscr{R}_{\ell-2,m}$ satisfying property $(\mathfrak{P}).$ Let $\mathcal{C}^{(\gamma_{\ell}+1)}$ be a $\Lambda_{\gamma_{\ell}+1}$-dimensional linear subcode of the torsion code $Tor_{1}(\mathscr{D}_{\ell-2}).$ The following hold.
@@
    $\displaystyle(2^{m})^{\sum\limits_{i=\ell}^{\gamma_{\ell}+\ell}\lambda_{i}\Lambda_{i-\ell+1}+\sum\limits_{j=\ell+1}^{\gamma_{\ell}+\ell}\lambda_{j}\Lambda_{j-\ell}+(\Lambda_{\gamma_{\ell}+\ell-1}+\Lambda_{\gamma_{\ell}+1})(n-\Lambda_{\gamma_{\ell}+\ell}-\Lambda_{\gamma_{\ell}+1})+\Lambda_{\gamma_{\ell}+1}^{2}+\Lambda_{\gamma_{\ell}+1}-Y_{\ell}}{\lambda_{\gamma_{\ell}+\ell}+n-\Lambda_{\gamma_{\ell}+\ell}-\Lambda_{\gamma_{\ell}+1}\brack\lambda_{\gamma_{\ell}+\ell}}_{2^{m}}\vskip-5.69054pt$
    distinct self-orthogonal codes $\mathscr{D}_{\ell}$ of type $\{\Lambda_{\gamma_{\ell}+1},\lambda_{\gamma_{\ell}+2},\ldots,\lambda_{\gamma_{\ell}+\ell}\}$ and length $n$ over $\mathscr{R}_{\ell,m}$ satisfying property $(\mathfrak{P})$ with $Tor_{1}(\mathscr{D}_{\ell})=\mathcal{C}^{(\gamma_{\ell}+1)}$ and $Tor_{i+1}(\mathscr{D}_{\ell})=Tor_{i}(\mathscr{D}_{\ell-2})$ for $1\leq i\leq\ell-2,$ where
[…]
@@
When $2\kappa>e$ and $e-\kappa+2-2\theta_{e}\leq\ell\leq\kappa-\mathrm{f}_{2\kappa-e}+1,$ the desired result follows immediately by working as in Proposition 3.3 and applying Lemma 4.1 of Yadav and Sharma [35]. On the other hand, when either $2\kappa\leq e$ and $\kappa+3\leq\ell\leq e$ or $2\kappa>e$ and $\kappa-\mathrm{f}_{2\kappa-e}+1<\ell\leq e,$ the desired result follows directly by proceeding as in Proposition 4.3 of Yadav and Sharma [34]. ∎
[…]
Figure 1: Flowchart illustrating the construction of a self-orthogonal code of type $\{\lambda_{1},\lambda_{2},\ldots,\lambda_{e}\}$ and length $n$ over $\mathscr{R}_{e,m}$ when $2\kappa\leq e$
[…]
Figure 2: Flowchart illustrating the construction of a self-orthogonal code of type $\{\lambda_{1},\lambda_{2},\ldots,\lambda_{e}\}$ and length $n$ over $\mathscr{R}_{e,m}$ when $2\kappa>e$
@@
We next assert that $\textbf{1}\notin Tor_{s-\kappa+\theta_{e}}(\mathscr{D}_{e})$ when $2\kappa\leq e,$ $n\equiv 4\pmod{8}$ and $m$ is odd. To prove this assertion, let us suppose that the code $\mathscr{D}_{e}$ has a generator matrix $\mathcal{G}_{e}$ as defined in (3.1). Further, corresponding to the code $\mathscr{D}_{e},$ let us consider a linear code $\mathscr{D}_{\ell}$ of type $\{\Lambda_{\gamma_{\ell}+1},\lambda_{\gamma_{\ell}+2},\ldots,\lambda_{\gamma_{\ell}+\ell}\}$ and length $n$ over $\mathscr{R}_{\ell,m}$ with a generator matrix $\mathcal{G}_{\ell}$ (as defined in (3.2)) for each integer $\ell$ satisfying $2\leq\ell\leq e-2$ and $\ell\equiv e\pmod{2}$. By Lemma 3.1, we see that $\mathscr{D}_{\ell}$ is a self-orthogonal code over $\mathscr{R}_{\ell,m}$ satisfying property $(\mathfrak{P}).$ In particular, $\mathscr{D}_{\kappa-1+\theta_{e}}$ is a self-orthogonal code of type $\{\Lambda_{s-\kappa_{1}+1},\lambda_{s-\kappa_{1}+2},\ldots,\lambda_{s+\kappa_{1}+\theta_{e}}\}$ over $\mathscr{R}_{\kappa-1+\theta_{e},m}$ satisfying property $(\mathfrak{P})$ and $\mathscr{D}_{\kappa+1+\theta_{e}}$ is a self-orthogonal code of type $\{\Lambda_{s-\kappa_{1}},\lambda_{s-\kappa_{1}+1},\ldots,\lambda_{s+\kappa_{1}+1+\theta_{e}}\}$ over $\mathscr{R}_{\kappa+1+\theta_{e},m}$ satisfying property $(\mathfrak{P}).$ Now, working as in Proposition 3.4, it is easy to see that $\textbf{1}\notin Tor_{s-\kappa+\theta_{e}}(\mathscr{D}_{e})$ when $2\kappa\leq e,$ $n\equiv 4\pmod{8}$ and $m$ is odd. On taking $\mathcal{C}^{(i)}=Tor_{i}(\mathscr{D}_{e})$ for $1\leq i\leq s+\theta_{e},$ the desired result follows. ∎